# An $\ell^{1}$-Plug-and-Play Approach for MPI Using a Zero Shot Denoiser with Evaluation on the 3D Open MPI Dataset

### v2

@@ 2.3 ZeroShot-PnP: The Proposed Plug-and-Play Approach with a Zero-Shot Denoiser
We remark here that if we consider no $\ell^{1}$-prior the same HQS yields the following iterative scheme, which can be regarded as a special case of Algorithm 1, which we will call ZeroShot-PnP in Section 3:
$\begin{split}u_{1}^{k+1}&=\arg\min_{u_{1}}\lVert f-Au_{1}\rVert_{2}^{2}+\mu_{k}\left\lVert u_{1}-u_{2}^{k}\right\rVert_{2}^{2}\\ u_{2}^{k+1}&=\mathrm{Denoiser}\left(u_{1}^{k+1}\,,\sqrt{\lambda/\mu_{k}}\right).\end{split}$ (12)
@@
Output: reconstructed volume $u$.
@@
4:     $u_{1}^{k+1}\leftarrow$ConjGrad$\left(A^{\ast}A+\mu_{k}\mathrm{Id}\,,A^{\ast}f+\mu_{k}\frac{u_{2}^{k}+u_{3}^{k}}{2}\right)$;
5:     $\sigma_{k}\leftarrow\sqrt{\mathrm{Var}(u_{1}^{k+1})}$;
@@
9:     $u_{2}^{k+1}\leftarrow$Slicewise-Denoise$\left(u_{1}^{k+1}\,,\sigma_{k}\right)$; $\triangleright$ ZeroShot-Denoiser
@@
11:     $\mu_{k+1}\leftarrow\lambda/\sigma_{k}^{2}$;
@@ 2.4 Parameter Selection Strategies
The ZeroShot-$\ell^{1}$-PnP approach proposed depends on the (hyper)parameters $\mu_{k}$, $\lambda$ and $\alpha$ as well as on the number of iterations $k$ of the scheme (stopping criterion). We remark that the $\sigma_{k}$ are not free parameters, but related via $\sigma_{k}=\sqrt{\lambda/\mu_{k}}$. Typically, in PnP approaches (e.g., [70]) the corresponding parameters are set fixed, in particular the noise related parameter $\sigma_{k}$ fed into the denoiser is set fixed for each iteration $k$ for a fixed number of iterations $n_{\mathrm{it}}$. We now propose more adaptive ways to choose the above parameters.
Choice of $\lambda$ and of the $\mu_{k}$
We start out from the relation $\sigma_{k}^{2}=\lambda/\mu_{k},$ noting that the $\sigma_{k}$ have an interpretation as (Gaussian) noise levels of the $u_{1}^{k}$. Using this interpretation, we let
$\tilde{\sigma}_{k}^{2}=\mathrm{Var}(u_{1}^{k})=\mathbb{E}\left[(u_{1}^{k}-\bar{u}_{1}^{k})^{2}\right],$ (13)
where $\bar{u}_{1}^{k}$ is the mean of the vector $u_{1}^{k}$. The value $\tilde{\sigma}_{k}^{2}$ is an upper bound on the noise level (and potentially can be seen as a very coarse estimate which empirically turns out to be sufficient for our purposes; another interpretation can be as a normalization). Then, we set the $\mu_{k}$ by
$\mu_{k}=\tilde{\mu}_{k}=\tilde{\lambda}/\tilde{\sigma}_{k}^{2}.$ (14)
To determine $\lambda=\tilde{\lambda},$ we use the relation $\tilde{\lambda}=\tilde{\mu}_{0}\tilde{\sigma}_{0}^{2},$ where we define $\tilde{\sigma}_{0}^{2}$ via Equation (13), and $\tilde{\mu}_{0}$ by the following observation: initializing $u_{2}^{0}=u_{3}^{0}=0$ in Algorithm 1, $u_{1}^{0}$ is the solution of the Tikhonov regularization problem (cf. Equation (9) and Equation (2) with $\mathcal{R}(u)=\lVert u\rVert_{2}^{2}$). Consequently, one can set the parameter $\mu_{0}$ to
$\tilde{\lambda}=\tilde{\mu}_{0}\tilde{\sigma}_{0}^{2},\qquad\tilde{\mu}_{0}=\tilde{\lambda}_{\mathrm{Tik}}$ (15)
@@
Finally, the algorithm is an iterative scheme that stops after $n_{\mathrm{it}}$ iterations and whose parameters are automatically updated at each iteration by the estimated parameter $\tilde{\sigma}_{k}$ in Equation (13). The only parameters that have to be provided are $\alpha$, $n_{\mathrm{it}}$ and the starting Tikhonov parameter $\mu_{0}=\tilde{\lambda}_{\mathrm{Tik}}$. We have validated $n_{\mathrm{it}}$ and $\mu_{0}$ on a simulated data set which we describe in Section 2.5.
Choice of $\alpha$
@@ 2.5 Hybrid Dataset for Parameter Validation
For the validation of the parameters of both the proposed methods and the baseline methods we have produced a hybrid MPI dataset. In more detail, we have produced concentrations $u_{\mathrm{GT}}$ on a $19\times 19\times 19$ grid that play the role of MPI phantoms. We have considered the system matrix $A$ provided in the Open MPI dataset and without preprocessing we have applied it to the ground truths $u_{\mathrm{GT}}$ to obtain a signal $f$ with realistic noise distribution (since $A$ is composed of real noisy delta scans) and added an additional Gaussian noise $\eta$ obtaining a dataset of scans of the form $f=Au_{\mathrm{GT}}+\eta$. The total dataset consists of 30 ground truths, generated to provide the same challenges for which the Open MPI dataset has been created, i.e., to test the reconstruction capabilities w.r.t. shape, different concentration levels and resolution. In particular, we have produced 10 cone-shaped phantoms, 10 graph-like phantoms and 10 additional phantoms composed of small dot-like regions with different concentration levels. The cones where produced with random sizes, shifts and orientations inside the grid. The graph like-phantoms were generated with the following procedure: a number of vertices $V$ between 4 and 6 is randomly chosen and $V-1$ distinct edges (pairs of vertices) are randomly selected; the voxels thorough which an edge passes or that contain a vertex get value $1$ whereas every other voxel stays zero. The resulting phantom is then blurred with a 3D normal Gaussian filter of variance 1 and thresholded to thicken the edges. Finally, the last set of phantoms is generated with the same procedure as for the graph-like phantoms but considering a randomly selected number of vertices between $6$ and $9$ but without edges. Each vertex gets a concentration level randomly selected between $0.05$ and $1$ and the Gaussian blurring and subsequent tresholding is applied to obtain regions bigger than a single voxel. All phantoms upon generation have maximum value 1 and have been rescaled to a maximum value $\beta\sim U(0.5\,,1.5)$.
We point out that this validation set is all we need to validate the parameters before testing on the Open MPI dataset and that no training set is necessary because no training is performed: we use a general purpose denoiser [70] pre-trained on natural images and we do not train and ad hoc denoiser.
@@ 3 Results
In this section we quantitatively and qualitatively evaluate both the ZeroShot-$\ell^{1}$-PnP in Algorithm 1 and the ZeroShot-PnP variant in Equation (12) in Section 2.3 on the OpenMPI dataset [39]. Qualitative comparision using the well established quality measures PSNR and SSIM is possible by inferring the ground truth from a CAD model; details are given in Section 3.1. We compare the ZeroShot-$\ell^{1}$-PnP and the ZeroShot-PnP schemes with the Deep Image Prior (DIP) and with the standard Tikhonov $\ell^{2}$-regularization as a baseline; moreover, we perform additional comparison with the PP-MPI algorithm [4] which uses a denoiser trained on MPI-friendly data. We show the performance of the PP-MPI denoiser when it is used in place of the ZeroShot-Denoiser in Algorihtm 1 (line 9).
Finally, we also show the performance of the PP-MPI denoiser in the ZeroShot-PnP algorithms. All these methods are described in Section 3.2. The evaluation is conducted in Section 3.3.
In particular, in Experiment 1 we compare the results obtained with the DIP, Tikhonov and PP-MPI methods on the 3D Open MPI Dataset using typical preprocessing and validated parameters. In Experiment 2 we focus more on the differences between the proposed ZeroShot-PnP method and the previous PP-MPI algorithm and the different performances of the denoisers employed. In Experiment 3 we test the ZeroShot-PnP with and without the $\ell^{1}$-prior with decreasing levels of preprocessing. Finally, in Experiment 4 we show that MPI reconstructions with the proposed methods are possible with minimal levels of preprocessing and without the employment of the SVD, which becomes more and more cumbersome to compute as the dimensionality of the problem increases.
The preprocessing was implemented both in Matlab and in Python while the ZeroShot-$\ell^{1}$-PnP and ZeroShot-PnP algorithms as well as the baseline algorithms for comparison were implemented in Python 3.9, using Numpy and PyTorch. The experiments were performed on a workstation with 13th Gen Intel(R) Core(TM) i9-13900KS, 128 GB of RAM, an NVIDIA RTX A6000 GPU and Windows 11 Pro.
@@ 3.1 Image Quality Measures
In order to quantitatively assess the reconstruction quality, the Open MPI dataset provides CAD models of the phantoms. This allows to compare the reconstructions with the corresponding ground truth implied by the CAD model. We use both Peak Signal-to-Noise Ratio (PSNR) and the Structural Similarity Index Measure (SSIM) [64] as measures.
@@
For the $\mathrm{SSIM}$ measure we employ the following definition and choice of parameters: given two images $f,g$, the SSIM is defined as $\mathrm{SSIM}(f,g)=\left[l(f,g)\right]^{\alpha}\cdot\left[c(f,g)\right]^{\beta}\cdot\left[s(f,g)\right]^{\gamma}$, where $l(f,g)=\frac{2\mu_{f}\mu_{g}+C_{1}}{\mu_{x}^{2}+\mu_{y}^{2}+C_{1}}$, $c(f,g)=\frac{2\sigma_{f}\sigma_{g}+C_{2}}{\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2}}$, $s(f,g)=\frac{\sigma_{fg}+C_{3}}{\sigma_{f}\sigma_{g}+C_{3}}$ are the luminance (l), contrast (c) and structure (s)  [64] and $\mu_{f}$ is the mean value, $\sigma_{f}$ the standard deviation of $f$; $\mu_{g}$ and $\sigma_{g}$ are analogously defined for $g$ and $\sigma_{fg}$ the covariance of $f$ and $g$. Following the implementation of [14], we set $\alpha=\beta=\gamma=1$ and $C_{1}=(0.01\cdot R)^{2}$, $C_{2}=(0.03\cdot R)^{2}$ and $C_{3}=0.5\cdot C_{2}$ where $R=100$, i.e., the concentration of the Delta Sample used during the calibration process. In what follows we always display $\mathrm{PSNR}_{\mathrm{max}}$ and $\mathrm{SSIM}_{\mathrm{max}}$ when showing the reconstruction of the Open MPI dataset.
3.2 Methods for Comparison
(Classical) Tikhonov Regularization
The inversion of the system matrix is performed considering the minimization problem in Equation  (2) with $\mathcal{R}(u)=\lVert u\rVert_{2}^{2}$; the minimizer is then obtained solving the associated Euler-Lagrange equations using the Conjugated Gradient (CG) method when the SVD is not available. When the SVD is indeed available, we use it for faster direct solutions of the inversion problems, it is however not necessary as we point out in Experiment 4.
Deep Image Prior (DIP)

### v3

@@ 2.2 The Preprocessing of the Data
If we consider no $\ell^{1}$-prior the same HQS yields the following iterative scheme, which can be regarded as a special case of Algorithm 1, which we call ZeroShot-PnP:
$\begin{split}u_{1}^{k+1}&=\arg\min_{u_{1}}\lVert f-Au_{1}\rVert_{2}^{2}+\mu_{k}\left\lVert u_{1}-u_{2}^{k}\right\rVert_{2}^{2}\\ u_{2}^{k+1}&=\mathrm{Denoiser}\left(\tilde{u}_{1}^{k+1}\,,\sqrt{\lambda/\mu_{k}}\right).\end{split}$ (12)
@@
For comparison we provide more details also on the PP-MPI denoiser [4]: the authors have first generated a dataset of MPI-like images obtained from time-of-flight MRA images from the public “ITKTubeTK - Bullitt - Healthy MR Dataset" by CASILab, by splitting the 95 healthy subjects into 77 training, 9 validation and 9 test sets; then, multiple $10\times 64\times 64$ 3D patches are cropped, thin-slab maximum-intensity projected along the first direction and finally downsampled to images of size $32\times 32$ and normalized to 1. Noise versions of these images have obtained by adding white Gaussian noise with standard deviation of 0.05. The PP-MPI denoiser uses a residual Dense network (RDN) [78] as its backbone architecture, adapted to have a compact set of parameters to match the small size of the MPI images. Concerning the number of network parameters the deep denoiser prior has $32,638,656$ parameters and the PP-MPI-Denoiser has $58,393$.
We conclude this section by observing that even though we have chosen to study and analyze the effect of the $\ell^{1}$-prior, the steps that led to the ZeroShot-$\ell^{1}$-PnP algorithm can be performed for any other prior $\mathcal{P}$ in place of the $\ell^{1}$ prior. In particular, the algorithm here proposed can be formulated to find minimizers of functionals of the form $\arg\min_{u}\lVert f-Au\rVert_{2}^{2}+\lambda\mathcal{R}(u)+\alpha\mathcal{P}(u)+\iota_{+}(u)$ for an additional convex prior $\mathcal{P}$; the convexity of $\mathcal{P}$ guarantees that its proximal mapping is well defined and can be directly substituted into Equation(11).
@@
Output: reconstructed volume $\tilde{u}_{\mathrm{rec}}$.
@@
4:     $u_{1}^{k+1}\leftarrow$ConjGrad$\left(A^{T}A+\mu_{k}\mathrm{Id}\,,A^{T}f+\mu_{k}\frac{u_{2}^{k}+u_{3}^{k}}{2}\right)$;
5:     $\sigma_{k+1}\leftarrow\sqrt{\mathrm{Var}(u_{1}^{k+1})}$;
@@
9:     $u_{2}^{k+1}\leftarrow$Slicewise-Denoise$\left(\tilde{u}_{1}^{k+1}\,,\sigma_{k+1}\right)$; $\triangleright$ ZeroShot-Denoiser
@@
11:     $\mu_{k+1}\leftarrow\lambda/\sigma_{k+1}^{2}$;
@@ 2.4 Parameter Selection Strategies
The ZeroShot-$\ell^{1}$-PnP approach proposed depends on the (hyper-)parameters $\mu_{k}$, $\lambda$ and $\alpha$ as well as on the number of iterations $k$ of the scheme (stopping criterion). We remark that the $\sigma_{k+1}$ are not free parameters, but related to $\lambda$ and $\mu_{k}$ via $\sigma_{k+1}=\sqrt{\lambda/\mu_{k}}$. Typically, in PnP approaches (e.g., [77]) the corresponding parameters are set fixed, in particular the noise related parameter $\sigma_{k}$ fed into the denoiser is set fixed for each iteration $k$ for a fixed number of iterations $n_{\mathrm{it}}$. We now propose more adaptive ways to choose the above parameters, as motivated by the continuation strategies in the previous section.
Choice of $\lambda$ and of the $\mu_{k}$.
We start out from the relation $\sigma_{k+1}^{2}=\lambda/\mu_{k},$ noting that the $\sigma_{k+1}$ have an interpretation as (Gaussian) noise levels of the $u_{1}^{k+1}$. Using this interpretation, we let
$\tilde{\sigma}_{k+1}^{2}=\mathrm{Var}(u_{1}^{k+1})=\mathbb{E}\left[(u_{1}^{k+1}-\bar{u}_{1}^{k+1})^{2}\right],$ (13)
where $\bar{u}_{1}^{k+1}$ is the mean of the vector $u_{1}^{k+1}$. The value $\tilde{\sigma}_{k+1}^{2}$ is an upper bound on the noise level (and potentially can be seen as a very coarse estimate which empirically turns out to be sufficient for our purposes; another interpretation can be as a normalization). Then, we set the $\mu_{k+1}$ by
$\mu_{k+1}=\tilde{\mu}_{k+1}=\tilde{\lambda}/\tilde{\sigma}_{k+1}^{2}.$ (14)
To determine $\lambda=\tilde{\lambda},$ we use the relation $\tilde{\lambda}=\tilde{\mu}_{0}\tilde{\sigma}_{0+1}^{2}=\tilde{\mu}_{0}\tilde{\sigma}_{1}^{2}$, where we define $\tilde{\sigma}_{1}^{2}$ via Equation (13), and $\tilde{\mu}_{0}$ by the following observation: initializing $u_{2}^{0}=u_{3}^{0}=0$ in Algorithm 1, $u_{1}^{1}$ is the solution of the Tikhonov regularization problem (cf. Equation (9) and Equation (2) with $\mathcal{R}(u)=\lVert u\rVert_{2}^{2}$). Consequently, one can set the parameter $\mu_{0}$ to
$\tilde{\lambda}=\tilde{\mu}_{0}\tilde{\sigma}_{1}^{2},\qquad\tilde{\mu}_{0}=\tilde{\lambda}_{\mathrm{Tik}}$ (15)
@@
Finally, the algorithm is an iterative scheme that stops after $n_{\mathrm{it}}$ iterations and whose parameters are automatically updated at each iteration by the estimated parameter $\tilde{\sigma}_{k+1}$ in Equation (13). The only parameters that have to be provided are $\alpha$, $n_{\mathrm{it}}$ and the starting Tikhonov parameter $\mu_{0}=\tilde{\lambda}_{\mathrm{Tik}}$. Even if $\mu_{0}$ can be formally interpreted as a Tikhonov parameter, in Section 3 we treat $\mu_{0}$ as a hyperparameter and we validate $n_{\mathrm{it}}$ and $\mu_{0}$ on a hybrid data set which we describe in Section 2.5.
Choice of $\alpha$.
@@ 2.5 Hybrid Dataset for Parameter Validation
For the validation of the parameters of both the proposed methods and the baseline methods we have produced a hybrid MPI dataset. In more detail, we have produced concentrations $u_{\mathrm{GT}}$ on a $19\times 19\times 19$ grid that play the role of MPI phantoms. We have considered the system matrix $A$ provided in the Open MPI dataset and without preprocessing we have applied it to the ground truths $u_{\mathrm{GT}}$ to obtain a signal $f$ with realistic noise distribution (since $A$ is composed of real noisy delta scans) and added an additional Gaussian noise $\eta$ obtaining a dataset of scans of the form $f=Au_{\mathrm{GT}}+\eta$. The total dataset consists of 30 ground truths, generated to provide the same challenges for which the Open MPI dataset has been created, i.e., to test the reconstruction capabilities w.r.t. shape, different concentration levels and resolution. In particular, we have produced 10 cone-shaped phantoms, 10 graph-like phantoms and 10 additional phantoms composed of small dot-like regions with different concentration levels. The cones where produced with random sizes, shifts and orientations inside the grid. The graph like-phantoms were generated with the following procedure: a number of vertices $V$ between 4 and 6 is randomly chosen and $V-1$ distinct edges (pairs of vertices) are randomly selected; the voxels thorough which an edge passes or that contain a vertex get value $1$ whereas every other voxel stays zero. The resulting phantom is then blurred with a 3D normal Gaussian filter of variance 1 and thresholded to thicken the edges. Finally, the last set of phantoms is generated with the same procedure as for the graph-like phantoms but considering a randomly selected number of vertices between $6$ and $9$ but without edges. Each vertex gets a concentration level randomly selected between $0.05$ and $1$ and the Gaussian blurring and subsequent tresholding is applied to obtain regions bigger than a single voxel. Because the delta probe of known and fixed maximum particle concentration $c_{\delta}=100$ mmol is used to acquire the system matrix, the outputs of the reconstruction methods will be rescaled reconstructions that shall be converted back to the absolute size via multiplication by $c_{\delta}$. All phantoms upon generation have maximum value 1 and have been rescaled to a maximum value $\beta\sim U(0.5\,,1.5)$. The range between 0.5 and 1.5 corresponds to phantoms with maximum concentration levels between 50 and 150 mmol, which is the correct order of magnitude for the phantoms in the Open MPI Dataset even if the Resolution Phantom does not fall precisely in this range. However, the quality of the reconstructions in Section 3 shows that such validation dataset works and that if it is generated with phantoms in the correct order of magnitude of the maximum concentration of the target phantom, the precise number does not affect too much the quality of the reconstruction.
We point out that this validation set is all we need to validate the parameters before testing on the Open MPI dataset and that no training set is necessary because no training is performed: we use a general purpose denoiser [77] pre-trained on natural images and we do not train an ad hoc denoiser.
@@ 3 Results
In this section we quantitatively and qualitatively evaluate both the ZeroShot-$\ell^{1}$-PnP in Algorithm 1 and the ZeroShot-PnP variant in Equation (12) in Section 2.3 on the OpenMPI dataset [43]. Qualitative comparision using the well established quality measures PSNR and SSIM is possible by inferring the ground truth from a CAD model; details are given in Section 3.1. We compare the ZeroShot-$\ell^{1}$-PnP and the ZeroShot-PnP schemes with the DIP, with the standard Tikhonov $\ell^{2}$-regularization and the ART method as a baseline; moreover, we perform additional comparison with the PP-MPI algorithm [4] which uses a denoiser trained on MPI-friendly data. We further investigate the performance of the PP-MPI denoiser when it is used in place of the ZeroShot-Denoiser in ZeroShot-$\ell^{1}$-PnP (Algorihtm 1, line 9). In each experiment, we employ the parameters obtained by evaluation on the hybrid dataset to reconstruct the phantoms in the Open MPI Dataset. All these methods are described in Section 3.2.
The evaluation is conducted in Section 3.3-3.6. In particular, in Experiment 1 (Section 3) we compare the results obtained with the DIP, Tikhonov, the ART and PP-MPI methods on the 3D Open MPI Dataset using typical preprocessing and the validated parameters. In Experiment 2 (Section 3.4) we focus more on the differences between the proposed ZeroShot-PnP method and the previous PP-MPI algorithm and the different performances of the denoisers employed. In Experiment 3 (Section 3.5) we test the ZeroShot-PnP with and without the $\ell^{1}$-prior with decreasing levels of preprocessing. Finally, in Experiment 4 (Section 3.6) we show that MPI reconstructions with the proposed methods are possible with most basic preprocessing (background removal and suppression of excitation crosstalk) and without the employment of the SVD, which becomes more and more cumbersome to compute as the dimensionality of the problem increases.
The preprocessing was implemented both in Matlab and in Python while the ZeroShot-$\ell^{1}$-PnP and ZeroShot-PnP algorithms as well as the baseline algorithms for comparison were implemented in Python 3.9, using Numpy and PyTorch. For the ART method as well as the PP-MPI method we have used the code made available by the authors of [4, 26] at https://github.com/icon-lab/PP-MPI. The experiments were performed on a workstation with 13th Gen Intel(R) Core(TM) i9-13900KS, 128 GB of RAM, an NVIDIA RTX A6000 GPU and Windows 11 Pro.
@@ 3.1 Image Quality Measures
In order to quantitatively assess the reconstruction quality, the Open MPI dataset provides CAD models of the phantoms. This allows to compare the reconstructions with the corresponding ground truth implied by the CAD model. We use both Peak Signal-to-Noise Ratio (PSNR) and the Structural Similarity Index Measure (SSIM) [71] as measures. We recall that that the PSNR of the reconstruction $f$ with respect to groundtruth $g$ is defined by
$\mathrm{PSNR}(f,g)=10\cdot\log_{10}\left(R^{2}/\mathrm{MSE}(f,g)\right)$ (16)
where $R$ is the maximum value of $f$ and $\mathrm{MSE}(f,g)$ is the Mean Square Error defined either as
$\mathrm{MSE}(f,g)=\frac{1}{MN}\sum_{j=1}^{M}\sum_{i=1}^{N}(f_{ij}-g_{ij})^{2}$ (17)
or as
$\mathrm{MSE}(f,g)=\frac{1}{LMN}\sum_{k=1}^{L}\sum_{j=1}^{M}\sum_{i=1}^{N}(f_{ijk}-g_{ijk})^{2},$ (18)
depending on if $f,g\in\mathbb{R}^{M\times N}$ are images or $f,g\in\mathbb{R}^{L\times M\times N}$ volumes.
For the $\mathrm{SSIM}$ measure we employ the following definition and choice of parameters: given two images $f,g$, the SSIM is defined as
$\mathrm{SSIM}(f,g)=\left[l(f,g)\right]^{\alpha}\cdot\left[c(f,g)\right]^{\beta}\cdot\left[s(f,g)\right]^{\gamma}$ (19)
where $l(f,g)=\frac{2\mu_{f}\mu_{g}+C_{1}}{\mu_{f}^{2}+\mu_{g}^{2}+C_{1}}$, $c(f,g)=\frac{2\sigma_{f}\sigma_{g}+C_{2}}{\sigma_{f}^{2}+\sigma_{g}^{2}+C_{2}}$, $s(f,g)=\frac{\sigma_{fg}+C_{3}}{\sigma_{f}\sigma_{g}+C_{3}}$ are the luminance (l), contrast (c) and structure (s)  [71] and $\mu_{f}$ is the mean value, $\sigma_{f}$ the standard deviation of $f$; $\mu_{g}$ and $\sigma_{g}$ are analogously defined for $g$ and $\sigma_{fg}$ the covariance of $f$ and $g$. Following the implementation of [15], we set $\alpha=\beta=\gamma=1$ and $C_{1}=(0.01\cdot R)^{2}$, $C_{2}=(0.03\cdot R)^{2}$ and $C_{3}=0.5\cdot C_{2}$ where $R=100$, i.e., the concentration of the Delta Sample used during the calibration process. All reconstructions are multiplied by the concentration level before the computation of the quality measures to maintain consistency with the choice of $R$ in the definition of the SSIM.
@@
In what follows we always display $\mathrm{PSNR}_{\mathrm{max}}$ and $\mathrm{SSIM}_{\mathrm{max}}$ when showing the reconstruction of the Open MPI dataset.
3.2 Methods for Comparison and Parameter Validation
(Classical) Tikhonov Regularization.
The inversion of the system matrix is performed considering the minimization problem in Equation  (2) with $\mathcal{R}(u)=\lVert u\rVert_{2}^{2}$; the minimizer is then obtained solving the associated Euler-Lagrange equations using the Conjugated Gradient (CG) method when the SVD is not available. When the SVD is indeed available, we use it for faster direct solutions of the inversion problems, it is however not necessary as we point out in Experiment 4. For the validation, we have chosen its parameter $\lambda_{\mathrm{Tik}}$ optimally in the sense that we have validated it performing a grid search and selecting that value for which the average $\mathrm{PSNR}$ on the hybrid dataset is maximal. The grid search proceeds as follows: we first perform reconstruction with parameters $10^{j}$ for $j=-6,-5,\dots,0,1,\cdots,18$ and consider $j^{\ast}$ the exponent that maximizes the PSNR; then we consider the values $k\cdot 10^{j^{\ast}-1}$ and $k\cdot 10^{j^{\ast}}$ with $k=1,\dots,9$ and again select the one maximizing the average PSNR.
Algebraic Reconstruction Technique (ART).
This is a method employed in MPI system-matrix-based reconstruction [34, 42]: the reconstruction is obtained solving the following constrained minimization problem $\hat{u}=\arg\min_{u\geq 0}\lVert f-Au\rVert_{2}^{2}+\lambda\lVert u\rVert_{2}^{2}$, using the ART method and incorporating a positivity constraint that is motivated by the fact that solution of the reconstruction a particle distributions. We have utilized the ART implementation provided in the PP-MPI author’s git repository [4, 26] available at https://github.com/icon-lab/PP-MPI. For the validation, we have performed the grid search for the regularization parameter $\lambda_{\mathrm{ART}}$ in the same fashion as for the Tikhonov method, performing 200 total iterations and considering selecting the regularization parameter and number of iterations $n_{\mathrm{it}}$ for which the PSNR is maximized after averaging on the hybrid validation set.
Deep Image Prior (DIP).